Hello, I need help with a couple of problems. I am totally lost!
y= x^2 + 2x and y = x^2 + 8x + 16 Any help would be greatly appreciated!

Answer by richwmiller(9143)

You said a couple of problems!
I see two equations.
Are they to be considered one problem or two?
y=x^2+2x
y=x^2+8x+16
x^2+2x=x^2+8x+16
0=6x+16
-16=6x
-16/6=x
-8/3=x
y=x^2+2x
y=(-8/3)^2+2*(-8/3)
y=64/9-16/3
y=64/6-48/9
y=16/9
(-8/3,16/9)